### v1

@@ Abstract
Scheduled for launch in 2030, the enhanced X-ray Timing and Polarization (eXTP) telescope is a Chinese space-based mission aimed at studying extreme conditions and phenomena in astrophysics. eXTP will feature three main payloads: Spectroscopy Focusing Arrays (SFAs), Polarimetry Focusing Arrays (PFAs), and a Wide-field Camera (W2C). This white paper outlines observatory science, incorporating key scientific advances and instrumental changes since the publication of the previous white paper [1]. We will discuss perspectives of eXTP on the research domains of flare stars, supernova remnants, pulsar wind nebulae, cataclysmic variables, X-ray binaries, ultraluminous X-ray sources, AGN, and pulsar-based positioning and timekeeping.
@@
In the new baseline design, the scientific payload of eXTP consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). The technical specifications and expected performance of these payloads can be found in White Paper 1 [2]
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 1.5 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of X-ray emitting compact objects at the shortest time scales. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device), to enhance imaging capabilities, which would supply strong upper limits to detect weak and extended sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of . Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
The PFA features three identical telescopes, with an angular resolution better than (HPD) in a FoV, and a total effective area of at (considering the detector efficiency). Polarization measurements are carried out by gas pixel detectors (GPDs) working at 2 – 8 keV with an expected energy resolution of 20% at 6 keV and a time resolution better than [7, 8, 9, 10, 11]. The instrument reaches an expected minimum detectable polarization (MDP) at confidence level () of about in for a milliCrab-like source.
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of (1 s exposure) across the 10–600 keV energy range, with an angular resolution of , a position accuracy of , and an energy resolution better than at 60 keV.
eXTP will be launched into a highly elliptical orbit with perigee and apogee altitudes of km and ,000 km, respectively. This high orbit allows continuous target monitoring observations, which are critical for studying variable X-ray sources that require uninterrupted long-duration (up to hr) observations.
@@
This white paper outlines the objectives of eXTP’s observatory science, incorporating key scientific advances since the previous white paper on these topics [1]. It will show how eXTP will address key questions in various research areas, from stars (Section 2), supernova remnants (SNRs) and pulsar wind nebulae (PWNe; Section 3), cataclysmic variables (CVs; Section 4), X-ray binaries (Section 5), ultraluminous X-ray sources (ULXs; Section 6), active galactic nuclei (AGN; Section 7), to pulsar-based timing and time-keeping (Section 8). In line with the current configuration of the mission, the revised goals emphasize the capabilities of SFAs and PFAs, while incorporating observational (e.g., by IXPE and NICER) and theoretical progress over the past few years. Moreover, this paper excludes the discussion of time-domain and multi-messenger science objectives originally presented in [1], as these topics now constitute a new science working group and will be presented in a separate white paper [6].
@@ 2 Flare stars
However, unlike the Sun, although there have been reports of CME candidates on some active late-type stars, there has not yet been a fully confirmed CME event (see, e.g., ref. [24]). Except for CME-induced effects like type II and moving type IV radio bursts, direct observational signatures of stellar coronae, such as Doppler shifts (or asymmetries) of emission lines and signals of coronal dimmings, are also considered feasible and promising (e.g., refs. [24, 25]). The soft X-ray band of eXTP can potentially capture signatures from stellar CMEs, and further constrain their properties, thereby facilitating more definitive detections of these events in the future.
In practice, long-term monitoring of stellar flares can be effectively conducted using smaller ground-based or space-based telescopes. Given that stellar flares can typically last for hours in the soft X-ray band (e.g., ref. [18]), there is a sufficient time window to trigger follow-up observations with eXTP once a flare has been detected. This strategy eliminates the need for eXTP to engage in continuous monitoring, thereby significantly reducing the observational cost for eXTP, while also ensuring a high likelihood of capturing a sufficient number of stellar flares and their accompanying phenomena.
Soft X-ray observations serve as a powerful tool for examining the characteristics of the stellar corona and transient eruptions such as flares and CMEs. By combining these observations with physical models, we can gain deeper insights into the mechanisms driving these phenomena and the role of magnetic fields in shaping stellar activity. Hereafter, we list a few important scientific questions about stars and stellar magnetic activities that will greatly benefit from the eXTP observations.
@@
Due to its large effective area, SFA can collect more X-ray photons from nearby stars than other X-ray telescopes with the same exposure time. Based on the differential emission measure (DEM) of Proxima Centauri at both the quiescent and flaring phases obtained by [26], we simulate 10-ks spectra of these two phases, respectively (Figure 2a). The flaring spectrum shows a flux times higher than the quiescent one. The spectral fit with a different DEM model can still distinguish these two phases by finding an excess of high-temperature components in the flaring phase.
The enhancement of high-temperature components (usually with a temperature of ) in the DEM during a flare (see, e.g., ref. [27]) primarily reflects localized heating in the corona, which causes the low-temperature plasma to rapidly heat up, shifting the emission measure from lower to higher temperatures. In addition, high-temperature emission may also result from chromospheric material that is heated and evaporates into the corona (e.g., ref. [28]).
@@
In addition, strong emission lines including Si XIII, Si XIV, S,XV, and even Fe,XXV can be observed in the current 10-ks spectrum, suggesting a probability of abundance measurement. Ar and Ca abundances can probably be further constrained with longer exposures.
@@
In history, simultaneous multi-wavelength observations play a key role in helping form the standard model of solar flares, in which magnetic reconnection releases magnetic energy and powers all kinds of flare radiation (e.g., ref. [30, 31]). Based on the scaling laws combining the statistical results of stellar flares at optical (e.g., refs. [15, 32, 33, 16, 34, 35, 36]) and soft X-ray (e.g., refs. [37, 18]) bands with solar flares, stellar flares are considered to have the same physical process as solar flares (e.g., ref. [38]). However, these deductions require validation by direct observations. Multi-wavelength monitoring of stellar flares offers a promising approach to this challenge. With its large effective area, the new-generation X-ray observatory eXTP has an outstanding photon-collecting ability, enabling spectroscopic observations with high temporal resolution. With a combination of continuous and simultaneous photometric and spectroscopic observations, eXTP will provide a strong constraint towards understanding the physical mechanisms driving stellar flares.
@@
Plasma ejection reduces the density of the source region, leading to a decreased emission of EUV and soft X-rays, which is known as coronal dimming (see, e.g., ref. [41]). With their excellent temporal resolution resulting from the large effective areas, the SFA instruments aboard eXTP make coronal dimming a promising method for stellar CME detection. For stars with a coronal temperature similar to that of the Sun, the characteristic lines of the corona are mostly at the EUV band. However, young or rapidly rotating stars tend to have higher coronal temperature and thus have stronger soft X-ray emissions (e.g., ref. [25]), making it possible to detect coronal dimmings in the X-ray band of eXTP. These stars are ideal targets for searching for CMEs via coronal dimming. The mass and velocity of CMEs are related to the amplitude and speed of the dimming of coronal emission, respectively (see, e.g., ref. [42]). The higher temporal resolution of SFA and SFA-I makes it possible to distinguish finer and weaker changes of flux in time, which enables eXTP to detect CME events less massive than former X-ray telescopes, and subsequently increases the detection rate within the same exposure time, since smaller CMEs are statistically more frequent (see, e.g., refs. [20, 18]). Such a higher detection rate will significantly enlarge the sample set of potential stellar CME candidates.
Moreover, the coronal dimming method for detecting CMEs requires an accurate diagnosis of the quiescent stellar corona, which serves as a basis for the calculation of the properties of potential CMEs. Thanks to the excellent photon-collecting ability of SFAs and SFA-I, the quiescent emission level from stellar coronae can be measured with shorter exposures. Consequently, the library of magnetically active stars with accurate quiescent coronal emission levels will also be significantly extended, thus establishing a robust basis for detecting stellar CMEs via the coronal dimming method.
@@ 3 Supernova remnants and pulsar wind nebulae
Supernovae result from energetic explosions of massive stars or white dwarfs. The ejected materials of supernovae interact with the interstellar medium and create supernova remnants (SNRs), each of which releases a huge amount of supernova mechanical energy (typically – erg). The core-collapse explosion of a massive star behind a neutron star or a black hole. In many cases, this may be a rapidly rotating, highly magnetized neutron star that powers a wind of charged particles, which interact with the surrounding magnetic fields and gas to form pulsar wind nebulae (PWNe). If the pulsar wind nebula (PWN) is young, it may still be embedded within an SNR. In such cases, the combined PWN/SNR system is classified as a subclass of SNRs, either as a plerion (or crab-like SNR) or, if a distinct shell is present, as a composite SNR [49].
SNRs and PWNe are among the brightest extended sources in the X-ray sky. They are important heating sources in galaxies and/or laboratories of high-energy phenomena. Through the diffusive shock acceleration mechanism, fast shocks of SNRs accelerate CR particles—relativistic protons, atomic nuclei, and electrons and, for PWNe, also positrons. CR electrons and positrons radiate synchrotron emission with a maximum emission at the characteristic energy keV [50], where and are the electron energy and the magnetic field component perpendicular to the motion of electrons. The characteristic cooling timescale of relativistic electrons due to synchrotron radiation is highly dependent on the electron/positron energy . As X-ray synchrotron typically comes from 10–100 TeV electrons, X-ray synchrotron radiation is only expected at or very near CR acceleration sites.
@@
Thanks to the excellent imaging polarization capabilities and large effective area of eXTP, the SNRs and PWNe will embrace new opportunities. Hereafter, we list a few important scientific questions of SNRs and PWNe that will greatly benefit from the eXTP observations.
@@
Despite a great deal of theoretical and observational progress, there are still many uncertainties regarding magnetic field amplification in and near SNRs. The compression of the interstellar medium alone only amplifies the magnetic fields by a factor of a few, so other processes are needed to explain the high amplification level. Several mechanisms have been proposed, such as amplification through CR-induced instabilities upstream of the shock [61] and due to turbulence dynamo processes downstream of the shock for SNRs evolving inside an air gap inhomogeneous gas environment [62, 63]. The upstream magnetic field turbulence is diminished by shock compression, which amplifies the tangential component. But downstream turbulence induced by density inhomogeneities —- which may in itself be induced by magnetic field turbulence [64]—may stretch again the magnetic fields in the radial directions [62].
@@
Polarization measurements provide a valuable way to study magnetic field turbulence and amplification since the polarization degree (PD) and polarization angle (PA) reflect the orderliness and orientation of the magnetic fields. Since the polarization direction can vary across the region, it is crucial to resolve the polarization structures on the smallest possible scales.
@@
IXPE has conducted a small survey of young SNRs with bright nonthermal X-ray emission, including Cas A [51], Tycho [53], SN 1006 [54, 55], RX J1713 [66], Vela Jr. [56]. See Table 1 for a summary of the results. These studies have detected X-ray polarization from all of these SNRs and demonstrated the unique value of imaging X-ray polarization in studying the magnetic fields in SNRs [67]. However, due to the small effective area of IXPE, none of the SNRs have their X-ray polarization resolved above the level at the resolution, despite the fact that IXPE observed near 1 Ms for each pointing of the SNRs. Moreover, the polarization detections are biased toward brighter regions or have been obtained by summing azimuthal regions, assuming a circular symmetry. The latter results in a depolarization if the polarization pattern deviates from pure circular symmetry. A case in point is Cas A for which the PD is %, but assuming a circular symmetric polarization pattern, while a more recent work provides 3–4 polarization detections showing that near the fainter forward shock region the PDs are 10%–26% [52].
@@
Therefore, limited by the small effective area or insufficient exposure time, IXPE has not fully applied its imaging capability for SNRs, while good resolution is important for preventing depolarization and constraining magnetic field properties. With a 5-fold larger effective area, eXTP can collect polarization signals more quickly, allowing more efficient measurements of the X-ray polarization distribution.
@@
eXTP will answer how magnetic turbulence and orientation vary across SNRs by measuring the distribution of X-ray polarization in young SNRs with a resolution . Cas A is an exceptionally suitable target for eXTP observations, as it represents not only one of the youngest and brightest SNRs in the X-ray band but also serves as a prominent prototype for comprehensive studies of SNR astrophysics. With 1 Ms observation with eXTP, we will obtain an X-ray polarization distribution in Cas A with unprecedented detail. The left panel of Figure 3 shows a simulated PFA map of the minimum detectable polarization at the 99% confidence level (MDP99) for SNR Cas A in the 3–6 keV band. With the PFA we can measure PDs without a brightness bias down to the PD=10% level at the rim and 2.5% at the reverse shock, using 30′′ pixel sizes. eXTP will show whether the PD and PA change between the forward shock and reverse shock and between different azimuthal angles, where shock velocities are different. eXTP will also tell whether the low PD in Cas A (significantly lower than other young SNRs with PD) is due to a depolarization effect. We also note that 1 Ms observation of a single object is long. However, Cas A will be observed with eXTP in multiple epochs as a calibration source. The observation time of Cas A will be accumulated over the years to increase the measurement statistics.
@@
IXPE observed seven brightest PWNe, that are Crab [71], Vela [72], MSH 15-52 [73], B0540-69 [74], G21.5-0.9, Kes 75, and 3C 58, the last three objects are still under study. These nebulae present a large diversity in polarization properties and magnetic field structures, likely associated with the acceleration mechanism, their ambient environment, and evolution stage. In general, these nebulae have a local high polarization degree that could reach up to 50%-60%, approaching the synchrotron limit. A spatially averaged PD is around 20% for Crab and B0540-69, and an unexpectedly high 45% for Vela.
Vela PWN is one of the most polarized and interesting nebula detected by IXPE [72]. The nebula mainly radiates in the X-ray and radio bands. The X-ray nebula is bright and highly structured in the central 1 arcmin square region, but rather weak and diffuse in the outskirt (2–4 arcmin). For now, IXPE only maps the X-ray polarization in the bright center. Considering the faintness of the diffuse emission, it is hard to obtain a robust result with 860 ks. With the same exposure performed by IXPE, eXTP can observe deeper and complete the missing part of the magnetic field to a larger area, as shown in Figure 4. This will provide a key to understanding the transport and diffusion of high-energy particles downstream. The result will have important implications for the efficiency of cosmic ray production.
Boomerang G106.6+2.9 is a middle-aged PWN with a highly polarized radio nebula similar to the Vela PWN. The nebula structure is believed to result from the interaction with the supernova reverse shock. [75] proposed that it may be highly polarized with a PD 45% in the X-ray band. However, the nebula is rather faint, making it not an ideal target for IXPE. With a larger effective area and higher sensitivity, eXTP could obtain a meaningful 5 PD detection using an exposure of 370 ks (assuming an integrated PD of 45%), which can help us better understand the nebula interacting with the reverse shock.
Many young pulsars have strong X-ray emission that could contaminate the magnetic field measurement of the surrounding PWN. A good time resolution of the eXTP is critical to help mitigate this issue, such that we can select only the off-pulse time period to perform phase-resolved analysis.

### v2

@@ Abstract
Scheduled for launch in 2030, the enhanced X-ray Timing and Polarization (eXTP) telescope is a Chinese space-based mission aimed at studying extreme conditions and phenomena in astrophysics. eXTP will feature three main payloads: Spectroscopy Focusing Array (SFA), Polarimetry Focusing Array (PFA), and a Wide-field Camera (W2C). This white paper outlines observatory science, incorporating key scientific advances and instrumental changes since the publication of the previous white paper [1]. We will discuss perspectives of eXTP on the research domains of flare stars, supernova remnants, pulsar wind nebulae, cataclysmic variables, X-ray binaries, ultraluminous X-ray sources, AGN, and pulsar-based positioning and timekeeping.
@@
In the new baseline design, the scientific payload of eXTP consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). The technical specifications and expected performance of these payloads can be found in [2]
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 6 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of X-ray emitting compact objects at the shortest time scales. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device), to enhance the capabilities of imaging and improve sensitivities to faint, point-like sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of . Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
The PFA features three identical telescopes, with an angular resolution better than (HPD) in a FoV, and a total effective area of at (considering the detector efficiency). Polarization measurements are carried out by gas pixel detectors (GPDs) working at 2 – 8 keV with an expected energy resolution of keV at 6 keV and a time resolution better than [7, 8, 9, 10, 11]. The instrument reaches an expected minimum detectable polarization (MDP) at confidence level () of about in for a milliCrab-like source.
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of (1 s exposure) across the 30–600 keV energy range, with an angular resolution of , a position accuracy of , and an energy resolution better than at 60 keV.
eXTP will be launched into a highly elliptical orbit with perigee and apogee altitudes of km and ,000 km, respectively, while a low-Earth orbit remains a viable alternative. This high orbit allows continuous target monitoring observations, which are critical for studying variable X-ray sources that require uninterrupted long-duration (up to hr) observations.
@@
This white paper outlines the objectives of eXTP’s observatory science, incorporating key scientific advances since the previous white paper on observatory science [1]. It will show how eXTP will address key questions in various research areas, from stars (Section 2), supernova remnants (SNRs) and pulsar wind nebulae (PWNe; Section 3), cataclysmic variables (CVs; Section 4), X-ray binaries (Section 5), ultraluminous X-ray sources (ULXs; Section 6), active galactic nuclei (AGN; Section 7), to pulsar-based timing and time-keeping (Section 8). In line with the current configuration of the mission, the revised goals emphasize the capabilities of the SFA and PFA, while incorporating observational (e.g., by IXPE and NICER) and theoretical progress over the past few years. Moreover, this paper excludes the discussion of time-domain and multi-messenger science objectives originally presented in [1], as these topics now constitute a new science working group and will be presented in a separate white paper [6].
@@ 2 Flare stars
However, unlike the Sun, although there have been reports of CME candidates on some active late-type stars, there has not yet been a fully confirmed CME event (see, e.g., ref. [24]). In addition to CME-induced effects like type II and moving type IV radio bursts, direct observational signatures of stellar coronae, such as Doppler shifts (or asymmetries) of emission lines and signals of coronal dimmings, are also considered feasible and promising (e.g., refs. [24, 25]). The soft X-ray band of eXTP can potentially capture signatures from stellar CMEs, and further constrain their properties, thereby facilitating more definitive detections of these events in the future.
In practice, long-term monitoring of stellar flares can be effectively conducted using smaller ground-based or space-based telescopes, and mainly focusing on certain magnetically active stars. Given that stellar flares can typically last for hours in the soft X-ray band (e.g., ref. [18]), there is a sufficient time window for triggering follow-up observations with eXTP once a flare has been detected. This strategy eliminates the need for eXTP to engage in continuous monitoring, thereby significantly reducing the observational cost for eXTP, while also ensuring a high likelihood of capturing a sufficient number of stellar flares and their accompanying phenomena. For each solar-type star, the estimated occurrence frequencies are , , and for stellar flares with soft X-ray energies of –, –, and , respectively [18]. Flares with such energy ranges from nearby stars () are easily detectable by eXTP according to the sensitivities of its payloads. eXTP will not respond to all stellar flare alerts but only observe those with important scientific values. Accordingly, the threshold for triggering eXTP follow-up observations can be set to a high level (e.g., corresponding soft X-ray energy ) for the investigation of high-energy flares and accompanying mass motions. Nearby M-dwarfs like Proxima Centauri also offer intriguing targets for X-ray follow-up observations due to their high magnetic activity and relevance to habitability studies.
Soft X-ray observations serve as a powerful tool for examining the characteristics of the stellar corona and transient eruptions such as flares and CMEs. By combining these observations with physical models, we can gain deeper insights into the mechanisms driving these phenomena and the role of magnetic fields in shaping stellar activity. Below, we list a few important scientific questions about stars and stellar magnetic activities that will greatly benefit from the eXTP observations.
@@
Due to its large effective area, SFA can collect more X-ray photons from nearby stars than other X-ray telescopes with the same exposure time. Based on the differential emission measure (DEM) of Proxima Centauri at both the quiescent and flaring phases obtained by [26], we simulate 10-ks spectra for each (Figure 2a). The flaring spectrum shows a flux times higher than the quiescent one. The spectral fit with another DEM model can still distinguish these two phases by finding an excess of high-temperature components in the flaring phase. The enhancement of high-temperature components (usually with a temperature of ) in the DEM during a flare (see, e.g., ref. [27]) primarily reflects localized heating in the corona, which causes the low-temperature plasma to rapidly heat up, shifting the emission measure from lower to higher temperatures. In addition, high-temperature emission may also result from chromospheric material that is heated and evaporates into the corona (e.g., ref. [28]).
@@
In addition, strong emission lines including Si XIII, Si XIV, S XV, and even Fe XXV can be observed in the current 10-ks spectrum, suggesting a probability of abundance measurement. Ar and Ca abundances can probably be further constrained with longer exposures.
@@
In history, simultaneous multi-wavelength observations play a key role in helping form the standard model of solar flares, in which magnetic reconnection releases magnetic energy and powers all kinds of flare radiation (e.g., ref. [30, 31]). Based on the scaling laws combining the statistical results of stellar flares at optical (e.g., refs. [15, 32, 33, 16, 34, 35, 36]) and soft X-ray (e.g., refs. [37, 18]) bands with solar flares, stellar flares are considered to have the same physical process as solar flares (e.g., ref. [38]). However, these deductions require validation by direct observations. Multi-wavelength monitoring of stellar flares offers a promising approach to this challenge. With its large effective area, the new-generation X-ray observatory eXTP has an outstanding photon-collecting ability that enables spectroscopic observations with high temporal resolution. With a combination of continuous and simultaneous photometric and spectroscopic observations, eXTP will provide a strong constraint towards understanding the physical mechanisms driving stellar flares.
@@
Plasma ejection reduces the density of the source region, leading to a decreased emission of EUV and soft X-rays, which is known as coronal dimming (see, e.g., ref. [41]). With their excellent temporal resolution resulting from the large effective areas, the SFA instruments aboard eXTP make coronal dimming a promising method for stellar CME detection. For stars with a coronal temperature similar to that of the Sun, the characteristic lines of the corona are mostly at the EUV band. However, young or rapidly rotating stars tend to have higher coronal temperature and thus have stronger soft X-ray emissions (e.g., ref. [25]), making it possible to detect coronal dimmings in the X-ray band of eXTP. Therefore, these stars are ideal targets for searching for CMEs via coronal dimming. The mass and velocity of CMEs are related to the amplitude and speed of the dimming of coronal emission, respectively (see, e.g., ref. [42]). The higher temporal resolution of SFA and SFA-I makes it possible to distinguish finer and weaker changes of flux in time, which enables eXTP to detect CME events less massive than former X-ray telescopes, and subsequently increases the detection rate within the same exposure time, since smaller CMEs are statistically more frequent (see, e.g., refs. [20, 18]). Such a higher detection rate will significantly enlarge the sample set of potential stellar CME candidates.
Moreover, the coronal dimming method for detecting CMEs requires an accurate diagnosis of the quiescent stellar corona, which serves as a basis for the calculation of the properties of potential CMEs. Thanks to the excellent photon-collecting ability of the SFA, the quiescent emission level from stellar coronae can be measured with shorter exposures. Consequently, the library of magnetically active stars with accurate quiescent coronal emission levels will also be significantly extended, thus establishing a robust basis for detecting stellar CMEs via the coronal dimming method.
@@ 3 Supernova remnants and pulsar wind nebulae
Supernovae result from energetic explosions of massive stars or white dwarfs. The ejected materials of supernovae interact with the interstellar medium and create supernova remnants (SNRs), each of which releases a huge amount of supernova mechanical energy (typically – erg). The core-collapse explosion of a massive star leaves behind a neutron star or a black hole. In some cases, the result can be a rapidly rotating, highly magnetized neutron star that powers a wind of charged particles, which interact with the surrounding magnetic fields and gas to form pulsar wind nebulae (PWNe). If the pulsar wind nebula (PWN) is young, it may still be embedded within an SNR. In such cases, the combined PWN/SNR system is classified as a subclass of SNRs, either as a plerion (or crab-like SNR) or, if a distinct shell is present, as a composite SNR [49].
SNRs and PWNe are among the brightest extended sources in the X-ray sky. They are important heating sources in galaxies and ideal laboratories for studying high-energy phenomena. Through the diffusive shock acceleration mechanism, fast shocks of SNRs accelerate CR particles—relativistic protons, atomic nuclei, and electrons and, for PWNe, also positrons. CR electrons and positrons radiate synchrotron emission with a maximum emission at the characteristic energy keV [50], where and are the electron energy and the magnetic field component perpendicular to the motion of electrons. The characteristic cooling timescale of relativistic electrons due to synchrotron radiation is highly dependent on the electron/positron energy . As X-ray synchrotron typically comes from 10–100 TeV electrons, X-ray synchrotron radiation is only expected at or very near CR acceleration sites.
@@
Thanks to the excellent imaging polarization capabilities and large effective area of eXTP, the SNRs and PWNe will embrace new opportunities. Below, we list a few important scientific questions of SNRs and PWNe that will greatly benefit from the eXTP observations.
@@
Despite a great deal of theoretical and observational progress, there are still many uncertainties regarding magnetic field amplification in and near SNRs. The compression of the interstellar medium alone only amplifies the magnetic fields by a factor of a few, so other processes are needed to explain the high amplification level. Several mechanisms have been proposed, such as amplification through CR-induced instabilities upstream of the shock [61] and due to turbulence dynamo processes downstream of the shock for SNRs evolving inside inhomogeneous gas environment [62, 63]. The upstream magnetic field turbulence is diminished by shock compression, which amplifies the tangential component. But downstream turbulence induced by density inhomogeneities —- which may in itself be induced by magnetic field turbulence [64]—may stretch again the magnetic fields in the radial directions [62].
@@
Polarization measurements provide a valuable way to study magnetic field turbulence and amplification since the polarization degree (PD) and polarization angle (PA) reflect the orderliness and orientation of the magnetic fields. Since the polarization direction can vary across the region, it is crucial to resolve the polarization structures on the smallest possible scales to avoid the depolarization.
@@
IXPE has conducted a small survey of young SNRs with bright nonthermal X-ray emission, including Cas A [51], Tycho [53], SN 1006 [54, 55], RX J1713 [66], Vela Jr. [56]. These observations have revealed a great variety in polarization properties among young SNRs: (1) those expanding into a tenuous medium show a higher PD, and (2) the three youngest SNRs exhibit radially oriented magnetic fields, whereas the two older ones exhibit tangential fields (see Table 1 for a summary). Furthermore, the X-ray polarization properties of SN 1006 are found to differ significantly from radio measurements, highlighting the distinct ability of X-ray polarimetry to probe amplied fields [55]. Together, these studies demonstrate the unique value of imaging X-ray polarimetry in investigating the magnetic fields of SNRs [67].
@@
Due to the small effective area of IXPE, none of the SNRs have their X-ray polarization resolved above the level at the resolution, despite the fact that IXPE observed near 1 Ms for each pointing of the SNRs. Moreover, the polarization detections are biased toward brighter regions or have been obtained by summing azimuthal regions, assuming a circular symmetry. The latter results in a depolarization if the polarization pattern deviates from pure circular symmetry. A case in point is Cas A for which the PD is %, but assuming a circular symmetric polarization pattern, while a more recent work provides 3–4 polarization detections showing that near the fainter forward shock region the PDs are 10%–26% [52].
@@
Therefore, limited by the small effective area or insufficient exposure time, IXPE has not fully utilized its imaging capability for SNRs. This is a significant limitation, as good resolution is important for preventing depolarization and constraining magnetic field properties. With a 5-fold larger effective area, eXTP can collect polarization signals more quickly, allowing more efficient measurements of the X-ray polarization distribution.
@@
eXTP will answer how magnetic turbulence and orientation vary across SNRs by measuring the distribution of X-ray polarization in young SNRs with a resolution . Cas A is an exceptionally suitable target for eXTP observations, as it represents not only one of the youngest and brightest SNRs in the X-ray band but also serves as a prominent prototype for comprehensive studies of SNR astrophysics. With 1 Ms observation with eXTP, we will obtain an X-ray polarization degree and angle distribution in Cas A with unprecedented detail. The left panel of Figure 3 shows a simulated PFA map of the minimum detectable polarization at the 99% confidence level (MDP99) for SNR Cas A in the 3–6 keV band. With the PFA we can measure PDs without a brightness bias down to the PD=10% level at the rim and 2.5% at the reverse shock, using 30′′ pixel sizes. eXTP will show whether the PD and PA change between the forward shock and reverse shock and between different azimuthal angles, where shock velocities are different. eXTP will also tell whether the low PD in Cas A (significantly lower than other young SNRs with PD) is due to a depolarization effect. We also note that 1 Ms observation of a single object is long. However, Cas A will be observed with eXTP in multiple epochs as a calibration source. The observation time of Cas A will be accumulated over the years to increase the measurement statistics.
@@
IXPE observed seven brightest PWNe, including Crab [71], Vela [72], MSH 15-52 [73], B0540-69 [74], G21.5-0.9, Kes 75, and 3C 58, the last three objects are still under study. These nebulae present a large diversity in polarization properties and magnetic field structures, likely associated with the acceleration mechanism, their ambient environment, and evolution stage. In general, these nebulae have a local high polarization degree that could reach up to 50%-60%, approaching the synchrotron limit. A spatially averaged PD is around 20% for Crab and B0540-69, and an unexpectedly high 45% for Vela.
Vela PWN is one of the most polarized and interesting nebula detected by IXPE [72]. The nebula mainly radiates in the X-ray and radio bands. The X-ray nebula is bright and highly structured in the central 1 arcmin square region, but rather weak and diffuse in the outskirt (2–4 arcmin). For now, IXPE only maps the X-ray polarization in the bright center. Due to the faintness of the diffuse emission, IXPE could not obtain a robust result with 860 ks. With the same exposure performed by IXPE, eXTP can observe deeper and complete the missing part of the magnetic field to a larger area, as shown in Figure 4. This will provide a key to understanding the transport and diffusion of high-energy particles downstream. The result will have important implications for the efficiency of cosmic ray production.
Boomerang G106.6+2.9 is a middle-aged PWN that has a highly polarized radio nebula similar to the Vela PWN. The nebula structure is believed to result from the interaction with the supernova reverse shock. [75] proposed that it may be highly polarized with a PD 45% in the X-ray band. However, the nebula is rather faint, making it not suboptimal target for IXPE. With a larger effective area and higher sensitivity, eXTP could obtain a meaningful 5 PD detection using an exposure of 370 ks (assuming an integrated PD of 45%), which can help us better understand the nebula interacting with the reverse shock.
Many young pulsars have strong X-ray emission that could contaminate the measurements of the magnetic fields in the surrounding PWN. A good time resolution of the eXTP is critical to help mitigate this issue, such that we can select only the off-pulse time period to perform phase-resolved analysis.